# Hybrid Pruning: In-Situ Compression of Self-Supervised Speech Models for Speaker Verification and Anti-Spoofing

### v1

@@ 1 Introduction
Self-supervised learning has advanced speech processing by providing versatile pretrained models for downstream tasks like automatic speech recognition (ASR), speaker verification (SV), and anti-spoofing [1, 2, 3]. Models like WavLM [4], trained on large-scale speech datasets, achieve strong performance across these tasks. However, their large size (e.g., 94 million parameters for WavLM-Base and over 300 million for WavLM-Large) makes deployment challenging in resource-constrained environments, such as mobile devices or edge systems.
One promising compression approach is structured pruning [5, 6, 7, 8], which removes entire components like attention heads or network layers to yield direct reductions in memory and latency. However, current pruning methods are often disconnected from the final application. They are typically applied separately either during pre-training [5, 6], agnostic to the downstream task, or as a post-hoc step after fine-tuning [7, 9]. This separation prevents the pruning process from being guided by the specific requirements of the end task, making it difficult to find an optimal compressed architecture.
@@
Overall, our contributions are the following: (1) We introduce a unified framework that integrates structured pruning directly into downstream fine-tuning, driven by task-specific classification loss, without the need for separate knowledge distillation or multi-stage pipelines. (2) We conduct a comprehensive evaluation across tasks, including SV (CNCeleb[10], VoxCeleb[11], anti-spoofing (ASVspoof5 [12], SpoofCeleb [13]), model scales (Base, Large), and dataset sizes (200 hours, 2000 hours). Our method achieves up to 70% sparsity with minimal performance loss on large datasets and improves results on low-resource datasets by mitigating overfitting. (3) We provide a detailed analysis of the resulting sparsity patterns, revealing task-dependent pruning signatures and showing how data scale influences architectural choices and representational shifts within the pruned model.
Fig. 1: The proposed Hybrid Pruning (HP) framework in comparison to prior work. (a) Our single-stage approach jointly optimizes for the downstream task with downstream back-end, such as speaker extractor back-end and structured sparsity, directly learning a pruned architecture. This avoids multi-stage pipelines like (b) task-agnostic pre-training pruning [5] and (c) post-hoc pruning [7]. Notably, our method (center detail) does not require the teacher-student knowledge distillation (right detail) common in other pruning techniques.
@@ 2 Related work
SSL in Speaker and Spoofing Detection. SSL representations have become foundational in both SV and anti-spoofing tasks. However, a widely used pipeline involves fine-tuning the entire large SSL model along with the downstream back-end [14, 15, 16, 17] such as ECAPA-TDNN [18] and MHFA [14], leading to high computational costs and a risk of overfitting, particularly when downstream data is limited [19]. Our work directly addresses this by proposing to simultaneously fine-tune and prune the SSL model. This approach not only simplifies the compression pipeline but also avoids reliance on a potentially suboptimal, pre-finetuned teacher model, allowing the student model to forge its own, more effective optimization path.
Structured Pruning of Speech Models. Structured pruning is an increasingly popular technique for compressing SSL models. Methods can be broadly categorized as task-agnostic and task-specific approaches. Task-agnostic method, such as DP-HuBERT [5, 6] as shown in Figure 1(b), integrates pruning into the SSL pre-training stage itself. Task-specific, as shown in Figure 1(c), apply pruning post-hoc to a model already fine-tuned for a task like ASR, often requiring knowledge distillation to recover performance [7, 20, 21]. Many existing downstream-focused methods follow a multi-stage process involving: (1) full fine-tuning, (2) pruning with distillation, and (3) a final recovery step. This paper proposes a hybrid approach to address this gap by integrating structured pruning directly into the fine-tuning stage, optimizing the model’s architecture and weights jointly in a single pass without dependence on a frozen teacher model.
@@ 3 The Hybrid Pruning Framework
We introduce Hybrid Pruning (HP), a framework that directly co-optimizes a model’s architecture and weights in a single optimization stage. Departing from multi-stage pipelines, HP simultaneously fine-tunes a model for a downstream task while structurally pruning it to a target sparsity. To achieve this with high precision on downstream tasks, we employ a gradient-based $L_{0}$ regularization method adapted from [22, 23].
@@ 3.1 Dynamic Architectural Adaptation via Stochastic Gates
To make this gating mechanism differentiable, we employ the Hard Concrete distribution [24, 25] to model each $z_{j}$. This is a continuous distribution that serves as a relaxation of the discrete Bernoulli distribution but has the critical property of being able to take an exact value of zero with non-zero probability. This allows the network to learn, via gradient descent, to completely “turn off” or partially prune certain components. The probability of a gate being zero is controlled by a learnable parameter $\alpha_{j}$, effectively making the model’s architecture itself a subject of the optimization process.
@@ 3.2 Joint Optimization for Performance and Sparsity
Task-Specific Loss ($\mathcal{L}_{\text{task}}$): The first term anchors the model to the downstream objective. For SV, we use the AAM Softmax loss [26] to learn discriminative embeddings. For anti-spoofing, we employ a standard Binary Cross-Entropy (BCE) loss. This term ensures that any architectural changes made by the pruning process are evaluated based on their impact on final task performance.
Pruning Regularizer ($\mathcal{R}_{\text{prune}}$): The second term drives the model towards a target sparsity level. Given the ideal regularizer is the non-differentiable $L_{0}$ norm, we adopt the tractable variational approximation from [22]. To precisely steer the model to a pre-defined target sparsity (i.e. the ratio of pruned parameters to the original model size), $t$, we use an augmented Lagrangian controller [23]. The regularizer is therefore the expected value of this controller over the distribution of the stochastic gates:
@@ 4.1 Setup
Datasets. We conduct a comprehensive evaluation of our proposed framework across two distinct tasks: SV and anti-spoofing. For SV, we utilize standard large-scale benchmarks. The models are trained on the development set of the VoxCeleb2 dataset[27, 11]. Then the performance is evaluated on the three official trials: Vox1-O, Vox1-E, and Vox1-H. To further assess the generalization capabilities of our method under different language and domain conditions, we also report performance on the CNCeleb corpus [28]. For anti-spoofing, we train on ASVspoof5 [12] and SpoofCeleb [13], and test on their respective official sets.
Implementation details. We utilize two scales of pre-trained SSL models: (1) The BASE models, consisting of a CNN encoder and 12 layers of Transformers with approximately 94M parameters. (2) The Large models, which include 24 layers of Transformers with approximately 316M parameters. For both SV and anti-spoofing tasks, we employ the same MHFA module with 32 attention heads, which accounts for a total of 1.2 million parameters [14]. During the training of our HP framework, the parameters of this back-end are jointly optimized along with the SSL model’s parameters and the pruning masks. For SV tasks, we employ the AAM-softmax loss with a margin of 0.2 and a scaling factor of 32. For the anti-spoofing task, we utilize a standard BCE loss. To ensure stable convergence during pruning, we implement a warm-up schedule for the target sparsity. The sparsity target is linearly increased from 0 to its final pre-defined value over the first 5 epochs of training.
@@
WavLM-SLIM (Best single system in ASVspoof5) [29] 5.16 0.149
@@ 4.2 Main Results on SV and Anti-Spoofing Benchmarks
For anti-spoofing, we use the standard ASVspoof5 benchmark, comparing our HP against two strong pruning baselines: a task-agnostic method (DP-HUBERT) and a post-hoc approach (Structured Pruning). As shown in Table 1, our HP method consistently and significantly outperforms both baselines at every sparsity level. Additionally, our method is the only one that improves performance over the unpruned baseline at 10% sparsity with 3.75% EER, which shows our method acts not only as a compression tool but as a powerful regularizer, as further analyzed in Sec 4.4. Furthermore, the FLOPs of our pruned model differ from the baselines at the same sparsity, which indicates distinct sub-architectures are being learned.
[…]

### v2

@@ 1 Introduction
Self-supervised learning has advanced speech processing by providing versatile pretrained models for downstream tasks like automatic speech recognition (ASR), speaker verification (SV), and anti-spoofing [hsu2021hubert, li2023parameter, wang22_odyssey]. Models like WavLM [chen2022wavlm], trained on large-scale speech datasets, achieve strong performance across these tasks. However, their large size (e.g., 94 million parameters for WavLM-Base and over 300 million for WavLM-Large) makes deployment challenging in resource-constrained environments, such as mobile devices or edge systems.
One promising compression approach is structured pruning [peng23c_interspeech, wang23da_interspeech, han2025efficient, ebrahimipour2025latency], which removes entire components like attention heads or network layers to yield direct reductions in memory and latency. However, current pruning methods are often disconnected from the final application. They are typically applied separately either during pre-training [peng23c_interspeech, wang23da_interspeech], agnostic to the downstream task, or as a post-hoc step after fine-tuning [han2025efficient, 10095780]. This separation prevents the pruning process from being guided by the specific requirements of the end task, making it difficult to find an optimal compressed architecture.
@@
Overall, our contributions are the following: (1) We introduce a unified framework that integrates structured pruning directly into downstream fine-tuning, driven by task-specific classification loss, without the need for separate knowledge distillation or multi-stage pipelines. (2) We conduct a comprehensive evaluation across tasks, including SV (CNCeleb[li2020cn], VoxCeleb[chung2018voxceleb2], anti-spoofing (ASVspoof5 [wang24_asvspoof], SpoofCeleb [jung2025spoofceleb]), model scales (Base, Large), and dataset sizes (200 hours, 2000 hours). Our method achieves up to 70% sparsity with minimal performance loss on large datasets and improves results on low-resource datasets by mitigating overfitting. (3) We provide a detailed analysis of the resulting sparsity patterns, revealing task-dependent pruning signatures and showing how data scale influences architectural choices and representational shifts within the pruned model.
Fig. 1: The proposed Hybrid Pruning (HP) framework in comparison to prior work. (a) Our single-stage approach jointly optimizes for the downstream task with downstream back-end, such as speaker extractor back-end and structured sparsity, directly learning a pruned architecture. This avoids multi-stage pipelines like (b) task-agnostic pre-training pruning [peng23c_interspeech] and (c) post-hoc pruning [han2025efficient]. Notably, our method (center detail) does not require the teacher-student knowledge distillation (right detail) common in other lightweight techniques [liao22_spsc].
@@ 2 Related work
SSL in Speaker and Spoofing Detection. SSL representations have become foundational in both SV and anti-spoofing tasks. However, a widely used pipeline involves fine-tuning the entire large SSL model along with the downstream back-end [peng2022attention, chen2022large, peng2025mhfa, rohdin24_asvspoof] such as ECAPA-TDNN [desplanques20_interspeech] and MHFA [peng2022attention], leading to high computational costs and a risk of overfitting, particularly when downstream data is limited [peng2023parameter]. Our work directly addresses this by proposing to simultaneously fine-tune and prune the SSL model. This approach not only simplifies the compression pipeline but also avoids reliance on a potentially suboptimal, pre-finetuned teacher model, allowing the student model to forge its own, more effective optimization path.
Structured Pruning of Speech Models. Structured pruning is an increasingly popular technique for compressing SSL models. Methods can be broadly categorized as task-agnostic and task-specific approaches. Task-agnostic method, such as DP-HuBERT [peng23c_interspeech, wang23da_interspeech] as shown in Figure 1(b), integrates pruning into the SSL pre-training stage itself. Task-specific, as shown in Figure 1(c), apply pruning post-hoc to a model already fine-tuned for a task like ASR, often requiring knowledge distillation to recover performance [han2025efficient, jiang23d_interspeech, han2025fine]. Many existing downstream-focused methods follow a multi-stage process involving: (1) full fine-tuning, (2) pruning with distillation, and (3) a final recovery step. This paper proposes a hybrid approach to address this gap by integrating structured pruning directly into the fine-tuning stage, optimizing the model’s architecture and weights jointly in a single pass without dependence on a frozen teacher model.
@@ 3 The Hybrid Pruning Framework
We introduce Hybrid Pruning (HP), a framework that directly co-optimizes a model’s architecture and weights in a single optimization stage. Departing from multi-stage pipelines, HP simultaneously fine-tunes a model for a downstream task while structurally pruning it to a target sparsity. To achieve this with high precision on downstream tasks, we employ a gradient-based $L_{0}$ regularization method adapted from [louizos2018learning, wang2020structured].
@@ 3.1 Dynamic Architectural Adaptation via Stochastic Gates
To make this gating mechanism differentiable, we employ the Hard Concrete distribution [maddison2016concrete, louizos2017learning] to model each $z_{j}$. This is a continuous distribution that serves as a relaxation of the discrete Bernoulli distribution but has the critical property of being able to take an exact value of zero with non-zero probability. This allows the network to learn, via gradient descent, to completely “turn off” or partially prune certain components. The probability of a gate being zero is controlled by a learnable parameter $\alpha_{j}$, effectively making the model’s architecture itself a subject of the optimization process.
@@ 3.2 Joint Optimization for Performance and Sparsity
Task-Specific Loss ($\mathcal{L}_{\text{task}}$): The first term anchors the model to the downstream objective. For SV, we use the AAM Softmax loss [deng2019arcface] to learn discriminative embeddings. For anti-spoofing, we employ a standard Binary Cross-Entropy (BCE) loss. This term ensures that any architectural changes made by the pruning process are evaluated based on their impact on final task performance.
Pruning Regularizer ($\mathcal{R}_{\text{prune}}$): The second term drives the model towards a target sparsity level. Given the ideal regularizer is the non-differentiable $L_{0}$ norm, we adopt the tractable variational approximation from [louizos2018learning]. To precisely steer the model to a pre-defined target sparsity (i.e. the ratio of pruned parameters to the original model size), $t$, we use an augmented Lagrangian controller [wang2020structured]. The regularizer is therefore the expected value of this controller over the distribution of the stochastic gates:
@@ 4.1 Setup
Datasets. We conduct a comprehensive evaluation of our proposed framework across two distinct tasks: SV and anti-spoofing. For SV, we utilize standard large-scale benchmarks. The models are trained on the development set of the VoxCeleb2 dataset[nagrani2017voxceleb, chung2018voxceleb2]. Then the performance is evaluated on the three official trials: Vox1-O, Vox1-E, and Vox1-H. To further assess the generalization capabilities of our method under different language and domain conditions, we also report performance on the CNCeleb corpus [fan2020cn]. For anti-spoofing, we train on ASVspoof5 [wang24_asvspoof] and SpoofCeleb [jung2025spoofceleb], and test on their respective official sets.
Implementation details. We utilize two scales of pre-trained SSL models: (1) The BASE models, consisting of a CNN encoder and 12 layers of Transformers with approximately 94M parameters. (2) The Large models, which include 24 layers of Transformers with approximately 316M parameters. For both SV and anti-spoofing tasks, we employ the same MHFA module with 32 attention heads, which accounts for a total of 1.2 million parameters [peng2022attention]. During the training of our HP framework, the parameters of this back-end are jointly optimized along with the SSL model’s parameters and the pruning masks. For SV tasks, we employ the AAM-softmax loss with a margin of 0.2 and a scaling factor of 32. For the anti-spoofing task, we utilize a standard BCE loss. To ensure stable convergence during pruning, we implement a warm-up schedule for the target sparsity. The sparsity target is linearly increased from 0 to its final pre-defined value over the first 5 epochs of training.
@@
WavLM-SLIM (Best single system in ASVspoof5) [zhu2024learn] 5.16 0.149
@@ 4.2 Main Results on SV and Anti-Spoofing Benchmarks
For anti-spoofing, we use the standard ASVspoof5 benchmark, comparing our HP against two strong pruning baselines: a task-agnostic method (DP-HUBERT) and a post-hoc approach (Structured Pruning). As shown in Table 1, our HP method consistently and significantly outperforms both baselines at every sparsity level. Additionally, our method is the only one that improves performance over the unpruned baseline at 10% sparsity with 3.70% EER, which shows our method acts not only as a compression tool but as a powerful regularizer, as further analyzed in Sec 4.4. Furthermore, the FLOPs of our pruned model differ from the baselines at the same sparsity, which indicates distinct sub-architectures are being learned.
As shown in Table 2, HP attains efficient inference while preserving accuracy for SV tasks. At 60% sparsity, WavLM Base+ reduces parameters from 95.6M to 38.9M and yields $2.2\times$ CPU and $2.0\times$ GPU inference speedup over the dense WavLM, with EERs of 0.70%, 0.78%, and 1.50% on Vox1-O/E/H, comparable to the dense model (0.70%, 0.69%, 1.40%).